OpenStudy (anonymous):

Perform each indicated operation. Simplify if possible. 4/2x + 9/3x

OpenStudy (anonymous):

nevermind, figured it out. find LCD, (3/3)4/2x + 9/3x(2/2) = 12/6x + 18/6x = 30/6x which simplified = 5/x